The perimeter of a rectangle is 48 m. If the width were doubled and the length were increased by 24 m, the perimeter would be 112 m. What are the length and width of the rectangle?

Answer by stanbon(75887)

The perimeter of a rectangle is 48 m. If the width were doubled and the length were increased by 24 m, the perimeter would be 112 m. What are the length and width of the rectangle?
-------
Perimeter = 2(length + width)
Original DATA
------
48 = 2(L + W)
L+W = 24 meters
----------------------
Modified DATA:
112 = 2(L+24 + 2W)
56 = L + 2W + 24
L + 2W = 32
-----
Solve the System::
L + W = 24
L + 2W = 32
--------------
Subtract and solve for "W:
W = 8 (original width)
L = 24-8 = 16 (original length)
--------------
Cheers,
Stan H.
--------------